OpenStudy (anonymous):

What are the vertices of the ellipse given by the equation 4x^2 + y^2 – 80x – 6y + 309 = 0?

hero (hero):

pk, solve this

hero (hero):

lol

hero (hero):

He ran away

OpenStudy (dumbcow):

write it in standard form \[\frac{(x-h)^{2}}{a^{2}} + \frac{(y-k)^{2}}{b^{2}} =1\] where vertices are center +- a

OpenStudy (anonymous):

dumbcow is correct, and @hero, I was helping coolaid earlier.

hero (hero):

You don't have to defend yourself pk

OpenStudy (dumbcow):

use completing the square rearrange equation to group x's and y's \[4x^{2}-80x +y^{2}-6y = -309\] \[4(x^{2}-20x) +y^{2}-6y = -309\] \[4(x^{2}-20x +100) +y^{2}-6y+9 = -309 +9 +4(100)\] \[4(x-10)^{2} + (y-3)^{2} = 100\] \[\frac{4(x-10)^{2}}{100} + \frac{(y-3)^{2}}{100} = 1\] \[\frac{(x-10)^{2}}{5^{2}} + \frac{(y-3)^{2}}{10^{2}} = 1\]

OpenStudy (dumbcow):

a is always larger a=10 b =5

OpenStudy (dumbcow):

center (10,3) a is under y so major axis is vertical vertices: (10,3+-10)